OpenStudy (anonymous):

Solve this system by substitution: Y = -5x + 8 -7x + 5y = -24 Please show work

Mehek (mehek14):

substitute y in the second equation so the second equation becomes -7x + 5(-5x + 8) = -24 do you understand so far?

Mehek (mehek14):

now solve it -7x + 5(-5x + 8) = -24 -7x + -25x + 40 = -24 (i multiplied 5 by -5x and got -25x and i multiplied 5 by 8 and got 40) -7x + -25x = -24 - 40 -7x + -25x = -64 -32x = -64 divide both sides by 2 x = 2

Mehek (mehek14):

did you understand? @mstrong_14